Lithium bromide as a aqueous solution with molar mass of 86.8g/mole and molality of 

1.35m and a density of 1.09g/mL. Calculate its molarity.


1
Expert's answer
2021-03-16T08:25:26-0400

"Density = \frac{mass}{volume}"


"Mass= density \u00d7 volume"


"Mass= 1.09\u00d71000= 1090g"


Mass of solution = mass of solute + mass of solvent

Mass of solution = (no. of moles of solute) × (molar mass of solute )+ mass of solvent


1800 = (1.35 × 86.8)+ mass of solvent


Mass of solvent = 1800 - 117.18 = 1682.82 g



"Molarity = \frac{(1.35×100)}{1682.82}=0.0802M"